Find an equation for the line: through (1, 8) and perpendicular to y = 5/4x + 4
Answer by anne(47)

ok
so perpendicular lines have negative reciprocals so our slope is -4/5
now we substiutute the points in the slope to find the y intercept

so
y = m x +b
8= -4/5(1) +b
8= -4/5 +b
8-4/5= -4/5+4/5 +b
40-4/5 =b
-36/5 =b

so our y intercept i -36/5

so our equation is y=-4/5x+-36/5

Did this help?
CHEERS